# The Unreasonable Effectiveness of Easy Training Data for Hard Tasks

### v1

@@ 4 Experiment Setup
Models. We conduct experiments with Llama-2 base models, for sizes 7b, 13b, and 70b Touvron et al. (2023b). In Appendix B, we show results for an RLHF version of Llama-2-70b, Qwen-72b Bai et al. (2023), and Mixtral-7x8b Jiang et al. (2024). Results are robust across models.
Data Hardness Stratification. To separate datasets into easy and hard data (with leftover data being medium data), we define easy/hard cutoffs as follows: for Question Num. Words, Answer Num. Chars, and MDL, we automatically define these values to be at the 30th and 70th percentiles of the variable range. Other variable cutoffs are defined manually: For ARC, Grade Level is easy (3-5), medium (6-7), hard (8); Difficulty Score is easy (1), medium (2), hard (3); Bloom Skill is easy (1-2), medium (3), hard (4-5). For MMLU, Grade Level is easy (high school) and hard (college) with no medium. For StrategyQA, Num. Reasoning Steps is easy (1-2), medium (3), hard (4-5). For GSM8k, Num. Reasoning Steps is easy (2-3), medium (4-5), hard (6-11).
Training Methods. We fit models to data with in-context learning (ICL; Brown et al., 2020), linear classifiers trained on frozen model hidden states Belinkov (2022), or QLoRA Dettmers et al. (2023). StrategyQA and GSM8k benefit heavily from utilizing chain-of-thought reasoning (CoT; Wei et al., 2022), so we primarily conduct experiments for these datasets with ICL+CoT and QLoRA+CoT (using reasoning chains from the datasets). See descriptions of each method below, with full detail in Appendix D.
Figure 3: Correlation heatmaps for ARC and GSM8k hardness measures.
Figure 4: Accuracy vs test data hardness across datasets (using Llama-2-70b with ICL). Data that humans find harder is also harder for LMs.
@@
    ICL: We leverage the in-context learning ability of LMs by placing “training” data into a prompt preceding each test question. We use $k=10$ in-context examples for ARC and MMLU and $k=8$ examples for StrategyQA and GSM8k (we see diminishing returns for larger $k$). When scoring multiple choice questions (no CoT), we get a model prediction by computing the answer probability for each answer choice given the test input and the prompt. When generating outputs with CoT, we greedily generate up to $t=100$ tokens for StrategyQA and $t=300$ tokens for GSM8k, then parse the string to get the predicted answer and compute exact-match accuracy against the label.
@@
    Linear Probing: We train a linear classifier on frozen LM hidden states. This is an effective method for performing multiple choice QA using LM representations Liu et al. (2023), and it does not require any finetuning of the underlying LM. For a given question, we compute one representation per answer choice by concatenating the question and answer choice and feeding it to the model. Then, we score each representation $z$ by applying the linear probe: $f(z;w)=w^{T}z$. The answer choice with the highest score is returned as the prediction. The probe weight $w$ is trained using SGD to minimize cross-entropy loss on a dataset of representations $Z=\{\{z_{i,j}\}_{j=1}^{j=|A|}\}_{i=1}^{N}$ derived from $N$ training datapoints with $|A|$ answer choices. Note we are not able to use this method for GSM8k as there are not multiple answer choices per datapoint.
@@
    QLoRA: Recently, low-rank adaptation of specific model parameters has enabled extremely parameter-efficient finetuning of LMs. To finetune our LMs, we execute QLoRA with the LoRA implementation from HuggingFace peft Mangrulkar et al. (2022) and the 8-bit AdamW from bitsandbytes Dettmers et al. (2022). We train the default layers for Llama-2 with rank $r=16$ adapters, $\alpha=32$, and dropout $p=0.1$. Model predictions are obtained in the same manner as for ICL, i.e., by scoring multiple choice options or generating $t=100/300$ tokens for StrategyQA/GSM8k.
@@
Statistical Testing. We perform experiments using 5 random seeds, controlling the training data selection (leaving remaining data for testing). To obtain confidence intervals and compute $p$-values for comparing results, we perform a block bootstrap to combine test and train data variance Efron and Tibshirani (1994). See Appendix E for further detail.
@@ 5 Experiments
Figure 5: Accuracy on hard test data vs train data hardness across datasets (using Llama-2-70b and ICL). Models recover 70-100% of the supervision gap (between Unsupervised and Hard) by training on Easy data.
5.1 RQ1: How Can We Measure Data Hardness? Do Different Approaches Agree?
We first explore the properties of our hardness measures for each dataset, to understand how they stand in comparison to one another.
Design. We first create correlation heatmaps for the hardness measures in our datasets, using a Spearman rank-order correlation Spearman (1987) between hardness values for each datapoint in a dataset. We show examples for ARC and GSM8k, which together provide complete coverage of our seven notions of data hardness (results for all datasets and hardness measures shown in Appendix Fig. 10). To validate that these measures are meaningful, we assess model performance across test data hardness levels, using Llama-2-70b and ICL with randomly sampled prompt examples.
Results. Based on Fig. 3, we find that our hardness measures are empirically very diverse. Off-diagonal correlations are usually fairly low, suggesting that these measures capture different possible aspects of datapoint hardness. [^3] Some of the ARC hardness correlations are low by design, as grade school exam questions are designed to vary in hardness within grade level (note the correlation of $\sim$0 between “Expert 1/2/3 Difficulty” and Grade Level). However, it appears that all of our hardness measures meaningfully capture some aspect of datapoint hardness, as model accuracy declines for harder test data for each of these measures (shown in Fig. 4), including for model-based hardness as measured by 7b-parameter models (shown in Appendix Fig. 11). Interestingly, we find that our model-based hardness measure is quite distinctive from our human-based measures, reaching a correlation with human hardness of at most 0.14 for ARC. This suggests that it is important to avoid relying only on model-based measures of hardness. Overall, we conclude that easy-to-hard generalization should be assessed with multiple notions of human hardness that are informed by cognitive science, since it appears that several meaningful measures of “hardness” do not strongly correlate with one another.
@@ 5.2 RQ2: Can We Do Well on Hard Data by Training on Easy Data?
Design. For each of our hardness measures, we test models on exclusively hard test data (according to that hardness measure), while varying whether they are trained on easy or hard data. We show results for Llama-2-70b using ICL. For results with additional hardness measures, see Appendix Fig. 19, and for other training methods, see Appendix Fig. 15. We focus on hard test data here, since that is the setting we are most interested in, but we provide results testing on the full data distribution in Appendix Fig. 13 and easy test data in Fig. 14.
Results. Surprisingly, Llama-2-70b shows comparable generalization to hard test data regardless of whether it is trained on easy or hard data (Fig. 5). In fact, across all six hardness measures, the Supervision Gap Recovered is between 70% and 100% (see SGR values with CIs and statistical significance tests in Appendix Table 2). Interestingly, for ARC and MMLU, there is no difference in easy vs. hard generalization using ICL. These results are robust across training methods (see Appendix Fig. 15). With QLoRA, the SGR also remains within 70%-100% for ARC, MMLU and StrategyQA. While GSM8k appears to exhibit worse easy-to-hard generalization, we note that easy-to-all generalization is actually equally good to hard-to-all generalization (see Fig. 13). Thus it seems like easy data provides surprisingly good supervision for LMs.
These results contrast notably with past work in curriculum learning and compositional generalization Bengio et al. (2009); Lake and Baroni (2018). This is likely because the models we consider, like Llama-2-70b, have learned much more during pretraining than models commonly used in work on curriculum learning and compositional generalization. So, it would seem that training these models on relatively small amounts of easy data successfully elicits the relevant task knowledge from the models in a way that is largely invariant to datapoint hardness.
@@
One implication of the results from Sec. 5.2 is that if easy data is almost as good as hard data, it could be better to collect and train on easy data, since hard data can be noisier and costlier to collect Sambasivan et al. (2021). Hence, we test the hypothesis that training on easy data outperforms hard data under two possible assumptions: (1) that one can collect more easy data than hard data given a fixed budget (i.e., time, money), and (2) that easy data is less noisily labeled than hard data.
Design. For (1) the budget assumption, we train linear probes on either easy or hard data using datasets of sizes in {10, 20, 40, 80, 160, 320}. We then show hard data test performance vs. training cost, assuming hard data costs twice as much as easy data to collect, meaning labeling 40 easy training points is equivalent in cost to labeling 20 hard training points. For (2) the noise assumption, we assume that easy data is mislabeled $p\%$ of the time, while hard data is mislabeled $2p\%$ of the time. Here, we measure test performance on hard data given different values of $p$. Note the 1:2 data collection cost ratio is almost exactly the ratio observed in MMLU-STEM-5, which contains 603 college level questions and 1143 high school questions, and a 1:2 labeling error ratio is plausible as well given expert human accuracy on datasets like MMLU (estimated at 89.8%) and GPQA (estimated at $\leq$72for difficult graduate level STEM questions) (Hendrycks et al., 2020; Rein et al., 2023). We use linear probes for this experiment given the expense of running 5 seeds of CoT experiments across different QLoRA training sizes for StrategyQA.
Figure 7: When hard train data has noisier labels than easy train data, training on easy data can give better hard test performance (shown for MMLU-STEM-5 using Llama-2-70b with a linear probe).
Results. We show results for the data budget assumption in Fig. 6. There is more often than not a clear advantage to fitting a model to easy data over hard data when the cost ratio between them is 1:2. The exceptions are for Bloom Skill and StrategyQA reasoning steps, where it appears that hard data does eventually meet or surpass the efficacy of easy training data. In the other three settings however, one can improve the model accuracy on hard test data by using easy rather than hard training data.
For the noise assumption, we draw a similar conclusion based on the results for MMLU in Fig. 7. It can be better to train on easy data rather than hard data, even when testing on hard test data. In this case, easy data becomes preferable when its labeling error rate is 10%, meaning the error rate for hard data is 20%. While these error rates are somewhat large, they are quite possible for difficult domain questions (Rein et al., 2023), suggesting that there are plausible settings where it is better to train on easy data than hard data due to label noise.
@@ 5.4 RQ4: Is Easy-To-Hard Generalization Consistent Across Model Scale and Train-Test Hardness Gap Size?
We attempt to answer two questions likely to be relevant as models become more capable: (1) how does the Supervision Gap Recovered change as models scale up, and (2) how does hard test performance change as the gap in train and test hardness grows? We are interested in these questions because in various settings AI performance may exceed human expert performance, and we want to know whether it will become more and more difficult to supervise models as this occurs.
Design. For question (1), we show easy-to-hard and hard-to-hard performance on MMLU for three different model sizes of Llama-2, namely 7b, 13b, and 70b. For question (2), we test models on hard MMLU data (college STEM questions), while training them on: high school MMLU data, 8th grade ARC data, 3rd grade ARC data, and no training data (unsupervised).
Results. First, we find that models show similar levels of Supervision Gap Recovered across scale (Fig. 8). SGR is 109% for the 7b model, 75% for the 13b, and 97% for the 70b, while the absolute difference between easy-to-hard and hard-to-hard performance is never more than 0.4 points. For our second question, we find that the difference between train and test hardness may have an effect on test performance (Fig. 9). Across methods, we see some decline in generalization once the gap between train and test hardness becomes sufficiently large, however 3rd grade supervision can be surprisingly effective for college STEM questions (e.g. SGR falls from 74% to 57% when training on 8th grade vs. 3rd grade questions using ICL). In an evaluation across our other hardness measures, easy training data is only marginally worse than medium training data (see Appendix Fig. 17). Together these results suggest that, while (1) the scalable oversight problem does not become harder as models scale up, (2) easy-to-hard performance may begin to decline when the gap between train and test hardness becomes sufficiently large.
Figure 9: Performance on hard test data may begin to decline when training on very easy data. Results shown for college-level STEM questions using Llama-2-70b (see other hardness measures in Appendix Fig. 17).
@@ 6 Discussion
Are Our Tasks Hard Enough to Provide Generalizable Easy-To-Hard Results? While humans can and have effectively labeled all of the datasets we use, the authors of this paper certainly could not have. Already, benchmark datasets for LMs require specialized domain expertise as benchmarks have increased in difficulty over time Clark et al. (2018); Hendrycks et al. (2020). The largest difficulty gap that we test in this paper is between 3rd grade and college level STEM questions (using ARC and MMLU). Concurrent work has called for work studying gaps as large as 3rd grade to 12th grade (described as “huge leaps in generalization”; Burns et al., 2023), and MMLU has been used in the past for scalable oversight experiments Bowman et al. (2022). Therefore, we see our results as relevant for future work that may operationalize “easy” and “hard” differently. One other interesting consideration here is that label noise in increasingly hard datasets may make experimental results less reliable as we test on harder and harder test sets, which could be an argument in favor of conducting experiments specifically on difficulty ranges like in MMLU where we have high trust that college-level test questions are labeled correctly.
How Do LMs Solve Hard Problems From As Few As Ten Easy Examples? Language models are known to be highly sample-efficient learners Brown et al. (2020), but our results demonstrate that they also efficiently learn to solve hard problems from easy data (recall that ICL achieves SGR scores of 70%-100% using no more than $k=10$ examples). How is this degree of supervision sufficient for recovering hard-to-hard performance? Evidently, training on even small amounts of easy data successfully elicits relevant knowledge from LMs in a way that is largely invariant to datapoint hardness. This could be because this kind of training encourages models to answer questions based on “truthfulness” representations of text, which should be invariant across domain and data hardness (see Azaria and Mitchell, 2023; Levinstein and Herrmann, 2023; Marks and Tegmark, 2023). We would point out that we do not interpret our results as models merely “learning the task format” or “learning the input/output space” as opposed to true generalization, because we are able to generalize to MMLU-STEM-5 college questions (a diverse mixture of physics, chemistry, biology, mathematics, and computer science questions) by using 3rd grade or 8th grade questions from ARC (Fig. 1), which come from a different dataset entirely. Hence it appears that training on easy data encourages models to give correct outputs for hard questions, though we note that the training signal is still underspecified in cases where SGR is less than 100%.
Why Not Use Test Questions About Knowledge From Beyond Pretraining Data Cutoffs? We want to emphasize that we are not trying to train models to answer questions for which there is no relevant information in their (pre)training data. Rather, we are trying to elicit knowledge from models that we suspect they may know, using fundamentally weak supervision. This is why hard-to-hard performance provides an upper bound for easy-to-hard generalization in our experiments (if a model simply cannot answer a question, the upper bound should reflect that). It would be interesting to consider test performance on “future” data given training data from before some cutoff, such that “easy” data is from the past and the “hard” training data is relevant “future” data.

### v2

@@ 4 Experiment Setup
Models. Apart from Fig. 1, we report results in the main paper on the Llama-2 70b base model Touvron et al. (2023b). In Appendix B, we show results for Llama-2 7b and 13b, an RLHF version of Llama-2-70b (“Llama-2-chat”), Qwen-72b Bai et al. (2023), and Mixtral-8x7b Jiang et al. (2024).
Data Hardness Stratification. To separate datasets into easy and hard data (with leftover data being medium data), we define easy/hard cutoffs as follows: for Question Num. Words, Answer Num. Chars, and MDL, we automatically define these values to be at the 30th and 70th percentiles of the variable range. Other variable cutoffs are defined manually: For ARC, Grade Level is easy (3-5), medium (6-7), hard (8); Difficulty Score is easy (1), medium (2), hard (3); Bloom Skill is easy (1-2), medium (3), hard (4-5). For MMLU, Grade Level is easy (high school) and hard (college) with no medium. For StrategyQA, Num. Reasoning Steps is easy (1-2), medium (3), hard (4-5). For GSM8k, Num. Reasoning Steps is easy (2-3), medium (4-5), hard (6-11). We show histograms for each hardness measure distribution in Appendix Fig. 20.
Finetuning Methods. We fit models to data with in-context learning (ICL; Brown et al., 2020), linear classifiers trained on frozen model hidden states Belinkov (2022), or QLoRA Dettmers et al. (2023). StrategyQA and GSM8k benefit heavily from utilizing chain-of-thought reasoning (CoT; Wei et al., 2022), so we primarily conduct experiments for these datasets with ICL+CoT and QLoRA+CoT (using reasoning chains from the datasets for supervision). See descriptions of each method below, with full detail in Appendix D.
@@
    ICL: For in-context learning (ICL), we use $k{=}10$ prompt examples for ARC and MMLU and $k{=}8$ examples for StrategyQA and GSM8k (we see diminishing returns for larger $k$). When scoring multiple choice questions (no CoT), we get a model prediction by computing the answer probability for each answer choice given the test input and the prompt. When generating outputs with CoT, we greedily generate up to $t=100$ tokens for StrategyQA and $t=300$ tokens for GSM8k. Accuracy is computed as exact match between predicted answer and label.
@@
    Linear Probing: We train a linear classifier on frozen LM hidden states. This is an effective method for performing multiple choice QA using LM representations Liu et al. (2023), and it does not require any finetuning of the underlying LM. For a given question, we compute one representation per answer choice by concatenating the question and answer choice as input and extracting the model’s final-token representation. Then, we score each representation $z$ by applying the linear probe: $f(z;w)=w^{T}z$. The answer choice with the highest score is returned as the prediction. The probe weight $w$ is trained using SGD to minimize cross-entropy loss on a dataset of representations $Z=\{\{z_{i,j}\}_{j=1}^{j=|A|}\}_{i=1}^{N}$ derived from $N$ training datapoints with $|A|$ answer choices.
@@
    QLoRA: To finetune our LMs, we execute QLoRA with the LoRA implementation from HuggingFace peft Mangrulkar et al. (2022) and the 8-bit AdamW from bitsandbytes Dettmers et al. (2022). We train the default layers for Llama-2 with rank $r=16$ adapters, $\alpha=32$, and dropout $p=0.1$. Model predictions are obtained in the same manner as for ICL, i.e., by scoring multiple choice options or generating $t=100/300$ tokens for StrategyQA/GSM8k.
@@
Training Size Controls. For all experiments with linear probing and QLoRA, we use $n=160$ train points. While we would prefer to use more finetuning data, the bottleneck we face is that fairly comparing easy-to-hard with hard-to-hard generalization requires both fixing the amount of finetuning data and leaving enough hard data left over for testing. Since we have as few as $n=603$ hard test points for MMLU, we have to limit finetuning data to $n=160$ points to leave enough test data for reasonably small confidence intervals.
Statistical Testing. We perform experiments using 5 random seeds, controlling the training data selection (leaving remaining data for testing). To obtain confidence intervals and $p$-values, we use block bootstrap sampling Efron and Tibshirani (1994). See Appendix E for further detail.
@@ 5 Experiments
Figure 3: Accuracy vs test data hardness across datasets (using Llama-2-70b with ICL). Data that humans find harder is also harder for LMs. Error bars are 95% CIs showing test sample variance.
Figure 4: Accuracy on hard test data vs train data hardness (using Llama-2-70b and ICL, averaged over 5 seeds). Models recover 70-100% of the supervision gap (between Unsupervised and Hard) when finetuned on Easy data.
5.1 RQ1: How Can We Measure Data Hardness?
We first explore properties of our hardness measures for each dataset. Here, we focus on hardness measures unique to each dataset, with full results across all measures shown in Appendix B.
Design. While our human hardness measures are direct measurements of data hardness, we validate that each measure is meaningful by assessing model performance across test hardness levels, using Llama-2-70b and ICL with randomly sampled prompts. We also create correlation heatmaps for hardness measures in our datasets, using a Spearman rank-order correlation Spearman (1987) between hardness values for each datapoint.
Results. It appears that all of our hardness measures meaningfully capture some aspect of datapoint hardness, as model accuracy declines for harder test data for each of these measures (Fig. 3), including for model-based hardness as measured by an ensemble of 7b-parameter models (Appendix Fig. 9). This also holds when finetuning with QLoRA (Appendix Fig. 10). Next, we find that our hardness measures are empirically very diverse (Appendix Fig. 8). Correlations between hardness measures are fairly low, suggesting that these measures capture different possible aspects of datapoint hardness. We conclude that easy-to-hard generalization should be assessed with multiple notions of datapoint hardness, since there are several different available measures and model performance declines for harder test data along each measure.
@@ 5.2 RQ2: Can We Do Well on Hard Data by Training on Easy Data?
Design. For each of our hardness measures, we test models on exclusively hard test data (according to that hardness measure), while varying whether they are finetuned on easy or hard data. [^3] We report test accuracy on the full data distribution and the easy test split in Appendix Figs. 11 and 12, respectively.
Results. Surprisingly, Llama-2-70b with ICL shows comparable generalization to hard test data regardless of whether it is fit to easy or hard data (Fig. 4). In fact, across all six hardness measures, the Supervision Gap Recovered is between 70% and 100%. These results are statistically significant, with CIs and $p$-values shown in Appendix Table 2. Interestingly, for ARC and MMLU, there is no difference in easy vs. hard generalization using ICL. Results are also robust across finetuning methods and additional hardness measures (Appendix Figs. 13, 18). With QLoRA, for example, the SGR remains within 70%-100% for ARC, MMLU and StrategyQA. While GSM8k appears to exhibit worse easy-to-hard generalization, we note that easy-to-all generalization is actually equally good to hard-to-all generalization (see Fig. 11). Thus it seems like easy data provides surprisingly good supervision for LMs.
These results contrast notably with past work in curriculum learning and compositional generalization Bengio et al. (2009); Lake and Baroni (2018). This is likely because models like Llama-2-70b have learned much more during pretraining than models commonly used in work on curriculum learning and compositional generalization. So, it would seem that finetuning these models on relatively small amounts of easy data successfully elicits the relevant task knowledge from the models in a way that is largely invariant to datapoint hardness.
One implication of the results from Sec. 5.2 is that if easy data is almost as good as hard data, it could be better to collect and fit to easy data, since hard data can be noisier and costlier to collect Sambasivan et al. (2021). Hence, we test the hypothesis that finetuning on easy data outperforms hard data under two possible assumptions: (1) that one can collect more easy data than hard data given a fixed budget (i.e., time, money), and (2) that easy data is less noisily labeled than hard data.
@@
Design. For (1) the budget assumption, we fit linear probes on either easy or hard data using datasets of sizes in {10, 20, 40, 80, 160, 320}. We then show hard data test performance vs. training cost, assuming hard data costs twice as much as easy data to collect, meaning labeling 40 easy training points is equivalent in cost to labeling 20 hard training points. For (2) the noise assumption, we assume that easy data is mislabeled $p\%$ of the time, while hard data is mislabeled $2p\%$ of the time. Here, we measure test performance on hard data given different values of $p$. Note the 1:2 data collection cost ratio is almost exactly the ratio observed in MMLU-STEM-5, which contains 603 college level questions and 1143 high school questions, and a 1:2 labeling error ratio is plausible as well given expert human accuracy on datasets like MMLU (estimated at 89.8%) and GPQA (estimated at $\leq$72for difficult graduate level STEM questions) (Hendrycks et al., 2021; Rein et al., 2023).
Results. We show results for the data budget assumption in Fig. 5. In terms of hard test accuracy, there is more often than not an advantage to fitting a model to easy data rather than hard data when the cost ratio between them is 1:2.
For the noise assumption, we draw a similar conclusion based on the results for MMLU in Fig. 6. For instance, easy data is preferable when its labeling error rate is 10% (or higher), meaning the error rate for hard data is 20% (or higher). Since high error rates are possible for difficult domain questions (Rein et al., 2023), there are plausible settings where it is better to finetune on easy data than hard data due to label noise.
@@ 5.4 RQ4: Is Easy-To-Hard Generalization Consistent Across Model Scale and Train-Test Hardness Gap Size?
Figure 6: When hard data has noisier labels than easy data, finetuning on easy data can give better hard test performance (shown for MMLU-STEM-5 using Llama-2-70b with a linear probe).
Figure 7: Performance on hard test data may begin to decline when finetuning on very easy data. Results shown for college-level STEM questions using Llama-2-70b (see other hardness measures, models in Appendix Figs. 16, 17).
We consider two questions likely to be relevant as models become more capable: (1) how does the Supervision Gap Recovered change as models scale in size, and (2) how does hard test performance change as the gap between train and test hardness grows? We are interested in these questions because in various settings AI performance may exceed human expert performance, and we want to know whether it will become more and more difficult to supervise models as this occurs.
Design. For question (1), we measure easy-to-hard and hard-to-hard performance on MMLU for Llama-2 at three different model sizes: 7b, 13b, and 70b. For question (2), we test models on hard MMLU data (college STEM questions), while finetuning them on even easier datasets than high school MMLU: 3rd and 8th grade ARC questions.
Results. First, we find that models show similar levels of Supervision Gap Recovered across sizes (Appendix Fig. 15). SGR is near 100% for model sizes between 7b and 70b. For our second question, we find that the difference between train and test hardness may have an effect on test performance (Fig. 7). Across methods, we see some decline in generalization once the gap between train and test hardness becomes sufficiently large. However, even 3rd grade supervision can be surprisingly effective for college STEM questions (e.g. SGR falls from 74% to 57% when fitting to 8th grade vs. 3rd grade questions using ICL). In an evaluation across our other hardness measures, easy training data is only marginally worse than medium training data (see Appendix Fig. 16). Together these results suggest that, while (1) easy supervision remains effective as models scale up, (2) easy-to-hard performance may begin to decline when the gap between train and test hardness becomes sufficiently large.
@@ 6 Discussion
Are Our Tasks Hard Enough to Provide Generalizable Easy-To-Hard Results? Benchmark datasets for LMs now require specialized domain expertise Hendrycks et al. (2021). The largest difficulty gap that we test in this paper is between 3rd grade and college level STEM questions (using ARC and MMLU). Concurrent work has called for work studying gaps as large as 3rd grade to 12th grade (described as “huge leaps in generalization”; Burns et al., 2023). Therefore, we see our results as relevant for future work that may operationalize “easy” and “hard” differently.
How Do LMs Solve Hard Problems From As Few As Ten Easy Examples? Our results suggest that finetuning on even small amounts of easy data successfully elicits relevant knowledge from LMs in a way that is largely invariant to datapoint hardness (we do not conclude that we are teaching LMs entirely new skills). This could be because this kind of finetuning encourages models to answer questions based on “truthfulness” representations of text, which should be invariant across domain and data hardness (see Marks and Tegmark, 2023). We emphasize that we do not interpret our results as models merely “learning the task format” as opposed to true generalization: we also fit models using ICL prompts that are trivially simple but match the task format for MMLU and StrategyQA, and find that model performance varies based on prompt data hardness and not simply prompt task format (see Appendix Fig. 19). Hence it appears that fitting to easy data encourages models to give correct outputs for hard questions.